if y^2=9; then y=?
Found 2 solutions by stanbon, shibili balakrishnan:
Answer by stanbon(48546)

if y^2=9; then y=?
y^2-9=0
(y-3)(y+3)= 0
y = 3 or y = -3
===================
Cheers,
Stan H.

Answer by shibili balakrishnan(3)

y^2>9
(y^2)^1/2>(9)^1/2
y>3